OpenStudy (anonymous):

Find the surface area for a cylinder with a radius of 4 and a height of 6. (use 3.14 for pi) 251.20 150.72 200.96 156.78

OpenStudy (anonymous):

OH Mhyy glob.. it's Jade.

OpenStudy (anonymous):

hah,hah, yeah, PLEASE HELP ME

OpenStudy (anonymous):

I am :)

OpenStudy (anonymous):

2 (pi r )* h

OpenStudy (zepp):



OpenStudy (anonymous):

2(4pi)*6

OpenStudy (zepp):

Ket-kat is here, go Ket-Kat :D

OpenStudy (anonymous):

Your answer will be: 150.8

OpenStudy (anonymous):

or 150.72 without rounding :D

OpenStudy (zepp):

@Ket-kat You made a little mistake in your formula, 2pi*r is the circonference, the area is pi *r^2

OpenStudy (zepp):

It should be 3.14 * 4^2 * 6

OpenStudy (anonymous):

Than, medals for you both i really needed this wanna help me with the last 2 lols

OpenStudy (anonymous):

How? That is the surface area equation?

OpenStudy (zepp):

Oh yeah, nvm, I misread the question :s

OpenStudy (anonymous):

301.44 isn't even an answer...

OpenStudy (zepp):

That's because you forgot to add the area of the bases :)

OpenStudy (anonymous):

No, I'm pretty sure I used the right equation lol. Anyway, there is your answer <3

OpenStudy (anonymous):

thanx, there are two more questions though, will you two stay

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Where are your other questions?

OpenStudy (zepp):

The answer should be 251.2 2*3.14*4+(3.14 * 4^2)*2 = 251.2

OpenStudy (anonymous):

Find the surface area of a cone with diamter of 13 and a slant height of 5 234.72 433.28 198.75 734.76

OpenStudy (zepp):

The surface area for a cone's formula is pi * r * s + pi * r^2

OpenStudy (anonymous):

pi(r)(s) + pi r^2

OpenStudy (zepp):

Radius = 13/2 = 6.5 pi * 6.5 * 5 + pi * 6.5^2 = 234.715

OpenStudy (anonymous):

Oh, Zepp is right

OpenStudy (anonymous):

Find the surface area for a prism with a perimeter of 24 area of a base 24 and a height of 20 480 528 48 480

OpenStudy (anonymous):

im still looking my computer is just slow

OpenStudy (anonymous):

It's all good :)